What is the pH of a buffer solution made of 50.0mL 1.00M acetic acid and 50.0 mL 1.0M sodium acetate?

What is the pH of 50mL of this buffer solution + 2.0mL 1.0M NaOH?

What is the pH of 50mL of this buffer solution + 2.0mL distilled water?
Answer #1

1) pKa of acetic acid = 4.74

50.0mL 1.00M acetic acid = 0.050 L * 1.00 mole / L = 0.050 mole.

and

50.0 mL 1.0 M sodium acetate = 0.050 L * 1.00 mole / L = 0.050 mole.

Henderson equation is

PH = pKa + log [salt] / [acid]

or

pH = 4.74 + log (0.050 / 0.050)

or

pH = 4.74

2) 2.0 mL of 1.0 M NaOH = 0.002 L * 1.0 mole / L = 0.002 mole.

PH = pKa + log [salt] / [acid]

or

pH = 4.74 + log [(0.050 + 0.002) / (0.050 - 0.002)]

pH = 4.77

3) when 2.0 mL distilled water added to 50 ml buffer solution then solution contains 0.050 mole of salt and 0.050 mole of acid.

hence

pH = 4.74